# Hierarchical Light Transformer Ensembles for Multimodal Trajectory Forecasting

### v1

@@ Abstract
Accurate trajectory forecasting is crucial for the performance of various systems, such as advanced driver-assistance systems and self-driving vehicles. These forecasts allow to anticipate events leading to collisions and, therefore, to mitigate them. Deep Neural Networks have excelled in motion forecasting, but issues like overconfidence and uncertainty quantification persist. Deep Ensembles address these concerns, yet applying them to multimodal distributions remains challenging. In this paper, we propose a novel approach named Hierarchical Light Transformer Ensembles (HLT-Ens), aimed at efficiently training an ensemble of Transformer architectures using a novel hierarchical loss function. HLT-Ens leverages grouped fully connected layers, inspired by grouped convolution techniques, to capture multimodal distributions, effectively. Through extensive experimentation, we demonstrate that HLT-Ens achieves state-of-the-art performance levels, offering a promising avenue for improving trajectory forecasting techniques.
Keywords:
Mixture Density Networks Efficient Ensembling Trajectory Forecasting
@@ 1 Introduction
Figure 1: Illustration of a hierarchical mixture (right) compared to a classical one (left) on a sample from the Argoverse 1 dataset. Ground truth is in dark blue in both figures, forecasts are colored dashed lines, and grey dashed lines are the centerlines of road lanes. On the right figure, we display the meta-modes (3, solid lines ending with a cross) inferred from the predictions, and their associated predictions (3, of the same color). The hierarchical structure notably enables efficient prediction compression by taking only the meta-modes.
@@
Indeed, achieving a longer prediction time horizon is challenging as the model has to account for complex interactions between varying numbers of traffic agents. As the uncertainty increases when forecasting further in time, predicting a single development becomes less relevant [23, 10, 15]. In this context, providing multiple outcomes (i.e., a set of forecasts) is more attractive but adds complexity to the task [15, 13, 40]. A prevalent approach involves the utilization of Mixture Density Networks (MDNs) [5] to capture the diverse modes (i.e., local maxima) of trajectory distributions efficiently [13, 33, 32, 49]. However, such methods face challenges in determining the appropriate number of modes, which is often ill-defined, and in ensuring that all the identified modes are not redundant. For instance, when approaching an intersection (see Fig. 1), there could be two distinct modes corresponding to turning right or proceeding straight. But, additional sub-modes may arise based on driver experiences or contextual factors. Furthermore, their industrial applications to safety-critical systems remain limited as they exhibit several shortcomings, such as overconfidence which is often typical in Deep Neural Networks (DNNs) [34, 14, 30]. Mitigating those issues, ensembles are attractive to enhance model accuracy [19] and strengthen resilience to uncertainty [21]. While they have been successfully applied to trajectory forecasting tasks [40, 42, 13, 49], their advantages come with increased computational overhead.
To address these challenges, we propose a novel solution based on hierarchical multi-modal density networks (Fig. 1, right) and efficient ensemble techniques. Our approach introduces a new loss function that captures the hierarchical structure of multi-modal predictions, allowing us to represent better the diverse behaviors exhibited by vehicles, particularly in complex scenarios. Additionally, building upon previous work [22], we adapt the concept of Packed-Ensembles, initially designed for Convolutional Neural Networks, to the realm of transformer-based models and develop a framework to train ensembles of Transformers efficiently without increasing computational complexity.
Our contributions extend beyond traditional ensemble learning techniques by focusing on both accuracy and uncertainty estimation in trajectory forecasting for autonomous driving systems. We demonstrate the effectiveness of our approach through extensive experimentation on trajectory forecasting datasets [8, 48], showcasing improved performance compared to existing methods. By leveraging hierarchical multi-modal networks and packed ensemble transformers, our method represents a significant advancement in trajectory forecasting, with potential applications in real-world autonomous driving scenarios.
In summary, our contributions are the following:
  • We introduce a novel density representation and associated loss function tailored for optimizing MDNs, leveraging a hierarchical structure to capture the nuanced nature of modes in trajectory forecasting.
@@ 2.1 Multimodal Trajectory Forecasting
When producing long-term forecasting, predicting a single forecast or an unimodal distribution is not enough. The future position distribution is more likely to follow a multimodal distribution [10, 15, 23, 13, 49]. For instance, without any prior information, the distribution density of vehicle future positions should be higher on the roads and lower outside. At the entry of an intersection, it might be relevant to estimate a trajectory for each possible direction. To that end, [26, 37] propose to approximate the modes of the distribution using several unimodal models, and [38, 13, 3, 9, 49], defines latent variables conditioning the decoder to generate a set of forecasts.
Yet, the predicted modes do not provide any information about the uncertainty around them. To better approximate aleatoric uncertainty (input data-dependent errors), researchers have turned to probabilistic approaches. In particular, generative models have been widely used. Leveraging implicit models, [23, 27, 36, 47, 29] define a sampling procedure to generate trajectories as if they were coming from the actual conditional distribution. However, there is no information on the number of samples required to cover the underlying distribution, limiting the application of such models on safety-critical applications. On the other hand, [40, 7, 10, 11, 31, 18] estimate explicitly the distribution density with Mixture Density Networks (MDNs) [5]. Although they enable flexibility to approximate uncertainties, such models are difficult to train as they suffer from training instabilities and mode collapse (i.e., several component probabilities tend to zero consistently across examples), for which we propose solutions.
@@ 2.2 Multiple Choice Learning
In multimodal trajectory forecasting, we are interested in learning several acceptable hypotheses rather than a single one. Introduced by [16], Multiple Choice Learning is a general framework for this task. It offers a means to create cooperation between multiple predictor heads. Eventually, each head becomes specialized on a subset of the data. Inspired by this approach, [24] adapted it to train true and implicit ensembles of DNNs. In particular, they introduced Stochastic Multiple Choice Learning (sMCL) [25] training scheme based on the Winner-Takes-All (WTA) loss. Although the dataset has only one realization for each input, it enables diversity among the heads. With the same objective, [35] use a shared architecture instead of several independent ones and demonstrate the benefits of MCL for multi-label tasks while having significantly fewer parameters than its ensemble counterpart. Interestingly, many recent works leveraging MDNs for multimodal trajectory forecasting employ WTA loss variants, easing the mixture optimization [49, 42, 28]. Ideally, the WTA loss leads to a Voronoi tesselation of the ground truth [35], which could be interpreted as a perfect k-means clustering. However, the k-means algorithm is known to be dependent on its initialization, which could lead to convergence problems. In addition, the WTA loss suffers from stability issues during training, referred to as hypothesis collapse, i.e., some hypotheses receive little to no gradients due to unfortunate initialization and better other hypotheses. To mitigate this issue, [35] relaxe the WTA formulation to ensure all hypotheses are slightly updated unconditionally to their performance. Yet, [28] showcase it results in non-winning hypotheses being slowly pushed to the equilibrium, creating a spurious mode in the distribution. Therefore, they introduce the Evolving WTA (EWTA) strategy, updating the top-$n$ winners instead of solely the best one and decreasing $n$ iteratively during training. To better estimate multimodal distributions, we introduce a hierarchy to capture diversity on various scales and help the optimization process.
@@ 3 Background
In this section, we present the mathematical formalism for this work and offer a brief background on training MDNs for trajectory forecasting. Appendix 0.A summarizes the main notations.
@@ 3.1 Trajectory Forecasting
Let us consider a dataset $\mathcal{D}=\{\mathbf{X}^{i}\}^{|\mathcal{D}|}_{i=1}$, where each instance $\mathbf{X}^{i}\in\mathbb{R}^{T\times A\times 2}$ represents the trajectories of $A$ traffic agents over $T$ timesteps: $\mathbf{X}^{i}=\mathbf{X}^{i}_{1:T}=(\mathbb{X}^{i}_{1},...,\mathbb{X}^{i}_{T})$. Here, $\mathbb{X}^{i}_{t}=(\mathbf{x}^{i}_{1}(t),...,\mathbf{x}^{i}_{A}(t))$ with $x^{i}_{j}(t)\in\mathbb{R}^{2}$ denotes the position in the 2-dimensional Euclidean space of agent $j\in\llbracket 1,A\rrbracket$ at time $t\in\mathbb{N}$ for the $i$th instance in $\mathcal{D}$.
@@ 3.2 Multiple Choice Learning
Within the Multiple Choice Learning framework, we aim to capture the complexity of phenomena better by replacing a single prediction with multiple ones. Consider a predictor providing $K$ estimates:
@@
A commonly used loss to optimize such models is the Winner-Takes-All (WTA) loss (or oracle loss), which computes the error $\ell$ for the closest prediction among the $K\in\mathbb{N}$ alternatives. As outlined by [35], this loss is defined it by:
@@
where we consider a Voronoi tesselation of the space $\mathbb{Y}=\cup^{K}_{k=1}\mathbb{Y}_{k}$ induced by $K$ unknown generators $g^{k}(X)$ and the loss $\ell$:
$\mathbb{Y}_{k}(X)=\{Y\in\mathbb{Y}\ \colon\ \ell(g^{k}(X),Y)<\ell(g^{j}(X),Y)\ \forall j\neq k\}$ (4)
As described in [35], Eq. 3 can be understood as the aggregation of the Voronoi cell losses generated by the predictions $f_{\theta}^{k}(X)$. To help understand the behavior of this loss, [35] assume $\ell$ is the $\ell_{2}-loss$ commonly used in regression problems and demonstrate Eq. 3 is minimal when the generators $g^{k}(X)$ and the predictors $f_{\theta}^{k}(X)$ are identical, and correspond to a centroidal Voronoi tesselation (cf. Theorem 1 of [35]). Hence, minimizing this loss can be seen as a clustering in the $\mathbb{Y}$ space conditioned on some $X\in\mathbb{X}$. Applied to trajectory forecasting, it produces diverse suitable forecasts for the same observation.
@@ 3.3 Mixture Density Networks
As detailed in [28], one can train MDNs using optimization strategies coming from the Multiple Choice Learning framework, by considering $f_{\theta}^{k}(X)$ to be the tuple $(\boldsymbol{\mu}_{\theta_{k}}(X),\mathbf{b}_{\theta_{k}}(X),\pi_{\theta_{k}}(X))$ and $\ell$ to be the Negative Log Likelihood (NLL).
@@ 4.1 Hierarchical Multimodal Distribution
Trajectory prediction involves generating multi-modal forecasts using DNNs. However, determining the appropriate number of modes (denoted as $K$) for training poses a challenge as it can vary due to multiple factors and lacks intuitive guidance. Additionally, the inherent multi-modal and uncertain nature of DNNs often leads to models with large variances, resulting in unreliable predictions. To address this challenge and improve prediction stability, we propose adapting the MDN framework to a Mixture of Mixtures, thereby introducing a novel architecture termed Hierarchical Mixture of Density Networks (HMDN).
@@
$P_{k^{\star}}(\mathcal{Y}|\mathcal{X}=X)=\sum_{k^{\prime}=1}^{K^{\prime}}\pi_{\theta_{k^{\star},k^{\prime}}}(X)\mbox{Laplace}(\boldsymbol{\mu}_{\theta_{k^{\star},k^{\prime}}}(X),\mathbf{b}_{\theta_{k^{\star},k^{\prime}}}(X))$ (6)
Consequently, the hierarchical multimodal distribution is equal to :
@@
where $\pi_{\theta_{k^{\star}}}(X)=\sum_{k^{\prime}\in K^{\prime}}\pi_{\theta_{k^{\star},k^{\prime}}}(X)$. It is noteworthy that we assume $K^{\prime}$ remains consistent across all meta-modes, although it could vary for each meta-mode. Therefore, the total number of modes is $K=K^{\star}\times K^{\prime}$. Furthermore, we define the mean and scale of the meta-modes respectively as: $\overline{\boldsymbol{\mu}}_{\theta_{k^{\star}}}(X)=\frac{1}{K^{\prime}}\sum_{k^{\prime}}\pi_{\theta_{k^{\star},k^{\prime}}}(X)\boldsymbol{\mu}_{\theta_{k^{\star},k^{\prime}}}(X)$ and $\overline{\mathbf{b}}_{\theta_{k}^{\star}}^{2}(X)=\frac{1}{2K^{\prime}}\sum_{k^{\prime}}\left(2\mathbf{b}_{\theta_{k^{\star},k^{\prime}}}^{2}(X)+\right.$ $\left.\boldsymbol{\mu}_{\theta_{k^{\star},k^{\prime}}}^{2}(X)\right)-\overline{\boldsymbol{\mu}}_{\theta_{k^{\star}}}^{2}(X)$. Introducing these meta-modes leads to more stable predictions, as demonstrated in Tab. 2. We analyze further the loss robustness to the variation of the mode number in Appendix 0.F. The following section describes the loss we developed to train an MDN with the concept of meta-modes.

### v3

@@ Abstract
Accurate trajectory forecasting is crucial for the performance of various systems, such as advanced driver-assistance systems and self-driving vehicles. These forecasts allow us to anticipate events that lead to collisions and, therefore, to mitigate them. Deep Neural Networks have excelled in motion forecasting, but overconfidence and weak uncertainty quantification persist. Deep Ensembles address these concerns, yet applying them to multimodal distributions remains challenging. In this paper, we propose a novel approach named Hierarchical Light Transformer Ensembles (HLT-Ens) aimed at efficiently training an ensemble of Transformer architectures using a novel hierarchical loss function. HLT-Ens leverages grouped fully connected layers, inspired by grouped convolution techniques, to capture multimodal distributions effectively. We demonstrate that HLT-Ens achieves state-of-the-art performance levels through extensive experimentation, offering a promising avenue for improving trajectory forecasting techniques. We make our code available at github.com/alafage/hlt-ens.
[…]
@@ 1 Introduction
Figure 1: Illustration of a hierarchical mixture (right) compared to a classical one (left) on a sample from the Argoverse 1 dataset. Ground truth is in dark blue in both figures, forecasts are colored dashed lines, and grey dashed lines are the centerlines of road lanes. On the right figure, we display the meta-modes (3 solid lines ending with a cross) inferred from the predictions and their associated predictions (3 of the same color). The hierarchical structure notably enables efficient prediction compression by taking only the meta-modes.
@@
Indeed, achieving a longer prediction time horizon is challenging as the model has to account for complex interactions between varying numbers of traffic agents. As the uncertainty increases when forecasting further in time, predicting a single development becomes less relevant [24, 10, 16]. In this context, providing multiple outcomes (i.e., a set of forecasts) is more attractive but adds complexity to the task [16, 14, 41]. A prevalent approach involves the utilization of Mixture Density Networks (MDNs) [5] to capture the diverse modes (i.e., local maxima) of trajectory distributions efficiently [14, 35, 34, 50]. However, such methods face challenges in determining the appropriate number of modes, often ill-defined, and ensuring that all the identified modes are distinct. For instance, when approaching an intersection (see Fig. 1), there could be two modes corresponding to turning right or proceeding straight. But, additional sub-modes may arise based on driver experiences or contextual factors. Furthermore, their industrial applications to safety-critical systems still need to be improved as they exhibit several shortcomings, such as overconfidence, which is often typical in Deep Neural Networks (DNNs) [36, 15, 32]. A research direction to mitigate these issues is ensembling. Ensembles usually enable the improvement of the model’s accuracy [20] and strengthen the resilience to uncertainty [22]. Although they have been successfully applied to trajectory forecasting tasks [41, 43, 14, 50], their advantages come with increased computational overhead at both the training and evaluation time.
We propose a novel solution to address these challenges based on hierarchical multi-modal density networks (Fig. 1, right) and efficient ensemble techniques. Our approach introduces a new loss function that presents a hierarchical structure in multi-modal predictions, allowing us to represent traffic agents’ diverse behaviors better. Additionally, based on previous work [23], we adapt the concept of Packed-Ensembles, initially designed for Convolutional Neural Networks, to the realm of transformer-based models and develop a framework to train ensembles of Transformers efficiently without increasing computational complexity.
We demonstrate the effectiveness of our approach through extensive experimentation on trajectory forecasting datasets [8, 49], showcasing improved performance compared to existing methods. In summary, our contributions are the following:
  • We introduce a novel density representation and associated loss function tailored for optimizing MDNs, leveraging a hierarchical structure to capture the nuanced nature of modes in trajectory forecasting and to embed spatial information in the prediction.
@@ 2.1 Multimodal Trajectory Forecasting
Predicting a single forecast or unimodal distribution is insufficient when producing long-term forecasting. The future position distribution will likely follow a multimodal distribution [10, 16, 24, 14, 50]. For instance, without any prior information, the distribution density of vehicle future positions should be higher on the roads and lower outside. At the entry of an intersection, it might be relevant to estimate a trajectory for each possible direction. To that end, [27, 39] approximate the modes of the distribution using several unimodal models, and [40, 14, 3, 9, 50], defines latent variables conditioning the decoder to generate a set of forecasts.
Yet, the predicted modes do not provide any information about the uncertainty around them. Thus, researchers have used probabilistic approaches to better approximate aleatoric uncertainty (input data-dependent errors). In particular, generative models have been widely used. Leveraging implicit models, [24, 28, 38, 48, 31] define a sampling procedure to generate trajectories as if they were coming from the actual conditional distribution. However, there is no information on the number of samples required to cover the underlying distribution, limiting the application of such models on safety-critical applications. On the other hand, [41, 7, 10, 11, 33, 19] estimate explicitly the distribution density with Mixture Density Networks (MDNs) [5]. Although they enable flexibility to approximate uncertainties, such models are difficult to train as they suffer from training instabilities and mode collapse (i.e., several component probabilities tend to zero consistently across examples), for which we propose solutions.
@@ 2.2 Multiple Choice Learning
In multimodal trajectory forecasting, we are interested in learning several acceptable hypotheses rather than one. Introduced by [17], Multiple Choice Learning is a general framework for this task. It offers a means to create cooperation between multiple predictor heads. Eventually, each head becomes specialized on a subset of the data. Inspired by this approach, [25] adapted it to train actual and implicit ensembles of DNNs. In particular, they introduced the Stochastic Multiple Choice Learning (sMCL) [26] training scheme based on the Winner-Takes-All (WTA) loss. Although the dataset has only one realization for each input, it enables diversity among the heads. With the same objective, [37] use a shared architecture instead of several independent ones and demonstrate the benefits of MCL for multi-label tasks while having significantly fewer parameters than its ensemble counterpart. Interestingly, many recent works leveraging MDNs for multimodal trajectory forecasting employ WTA loss variants, easing the mixture optimization [50, 43, 29]. Ideally, the WTA loss leads to a Voronoi tesselation of the ground truth [37], which could be interpreted as a perfect k-means clustering. However, similarly to the k-means algorithm, its performance depends on the hypotheses’ initialization. Indeed, the WTA loss is known to suffer from stability issues during training, referred to as hypothesis collapse, i.e., some hypotheses receive little to no gradients due to unfortunate initialization and better other hypotheses. To mitigate this issue, [37] relaxes the WTA formulation ($\varepsilon$-WTA) to ensure all hypotheses are slightly updated unconditionally to their performance. Yet, [29] showcase it results in non-winning hypotheses being slowly pushed to the equilibrium, creating a spurious mode in the distribution. Therefore, they introduce the Evolving WTA (EWTA) strategy, updating the top-$n$ winners instead of solely the best one and decreasing $n$ iteratively during training. To better estimate multimodal distributions, we introduce a hierarchy to capture diversity on various scales and help the optimization process.
@@ 2.3 Efficient Ensembling
Ensembling has been applied to trajectory forecasting problems. Yet the original formulation of Deep Ensembles for regression tasks is not suited to the Multiple Choice Learning framework where we want to retain diversity. Instead, we are interested in finding an efficient method to use sets of predictions to create a single set. In [30], the authors present the Robust Imitative Planning (RIP) algorithm, which keeps the hypotheses having the higher likelihoods on the estimated distributions from the ensemble. The issue is that it heavily relies on the quality of the predicted distributions and does not ensure any diversity. For better coverage, clustering approaches such as the Expectation-Maximization [41] or the K-Means [43, 50, 34] algorithms are attractive. In our experiments, we provide a benchmark of these ensembling strategies.
@@ 3 Background
In this section, we present the mathematical formalism for this work and offer a brief background on training MDNs for trajectory forecasting. Appendix A summarizes the main notations.
@@ 3.1 Trajectory Forecasting
Let us consider a dataset $\mathcal{D}=\{\mathbf{X}^{i}\}^{|\mathcal{D}|}_{i=1}$, where each instance $\mathbf{X}^{i}\in\mathbb{R}^{T\times A\times 2}$ represents the trajectories of $A$ traffic agents over $T$ timesteps: $\mathbf{X}^{i}=\mathbf{X}^{i}_{1:T}=(\mathbb{X}^{i}_{1},...,\mathbb{X}^{i}_{T})$. Here, $\mathbb{X}^{i}_{t}=(\mathbf{x}^{i}_{1}(t),...,\mathbf{x}^{i}_{A}(t))$ with $x^{i}_{a}(t)\in\mathbb{R}^{2}$ denotes the position in the 2-dimensional Euclidean space of agent $a\in\llbracket 1,A\rrbracket$ at time $t\in\mathbb{N}$ for the $i$th instance in $\mathcal{D}$.
@@ 3.2 Multiple Choice Learning
Within the Multiple Choice Learning framework, we aim to better capture phenomena’ complexity by replacing a single prediction with multiple ones. Consider a predictor providing $K$ estimates:
@@
A commonly used loss to optimize such models is the Winner-Takes-All (WTA) loss (or oracle loss), which computes the error $\ell$ for the closest prediction among the $K\in\mathbb{N}$ alternatives. As outlined by [37], this loss is defined by:
@@
where we consider a Voronoi tesselation of the space $\mathbb{Y}=\cup^{K}_{k=1}\mathbb{Y}^{k}$ induced by $K$ unknown generators $g^{k}(X)$ and the loss $\ell$:
$\mathbb{Y}^{k}(X)=\{Y\in\mathbb{Y}\ \colon\ \ell(g^{k}(X),Y)<\ell(g^{j}(X),Y)\ \forall j\neq k\}$ (4)
As described in [37], Eq. 3 can be understood as the aggregation of the Voronoi cell losses generated by the predictions $f_{\theta}^{k}(X)$. To help understand the behavior of this loss, [37] assume $\ell$ is the $\ell_{2}-loss$ commonly used in regression problems and demonstrate Eq. 3 is minimal when the generators $g^{k}(X)$ and the predictors $f_{\theta}^{k}(X)$ are identical, and correspond to a centroidal Voronoi tesselation (cf. Theorem 1 of [37]). Hence, minimizing this loss can be seen as a clustering in the $\mathbb{Y}$ space conditioned on some $X\in\mathbb{X}$. Applied to trajectory forecasting, it produces diverse, suitable forecasts for the same observation.
@@ 3.3 Mixture Density Networks
As detailed in [29], one can train MDNs using optimization strategies coming from the Multiple Choice Learning framework by considering $f_{\theta}^{k}(X)$ to be the tuple $(\boldsymbol{\mu}_{\theta_{k}}(X),\mathbf{b}_{\theta_{k}}(X),\pi_{\theta_{k}}(X))$ and $\ell$ to be the Negative Log Likelihood (NLL).
@@ 4.1 Hierarchical Multimodal Distribution
Trajectory prediction involves generating multi-modal forecasts using DNNs. However, determining the appropriate number of modes (denoted as $K$) for training poses a challenge as it can vary due to multiple factors and lacks intuitive guidance. Additionally, the inherent multi-modal and uncertain nature of DNNs often leads to models with significant variances, resulting in unreliable predictions. To address this challenge and improve prediction stability, we propose adapting the MDN framework to a Mixture of Mixtures, thereby introducing a novel architecture: Hierarchical Mixture of Density Networks (HMDN).
@@
$P_{k^{\star}}(\mathcal{Y}|\mathcal{X}=X)=\sum_{k’=1}^{K’}\pi_{\theta_{k^{\star},k^{\prime}}}(X)\mbox{Laplace}(\boldsymbol{\mu}_{\theta_{k^{\star},k^{\prime}}}(X),\mathbf{b}_{\theta_{k^{\star},k^{\prime}}}(X))$
We note $\theta_{k^{\star},k^{\prime}}$ the parameters of the $k^{\prime}$-th component of the $k^{\star}$-th meta-mode. Consequently, the hierarchical multimodal distribution is equal to :
@@
where $\pi_{\theta_{k^{\star}}}(X)=\sum_{k^{\prime}\in K^{\prime}}\pi_{\theta_{k^{\star},k^{\prime}}}(X)$. We assume $K^{\prime}$ remains consistent across all meta-modes, although it could vary for each. Therefore, the total number of modes is $K=K^{\star}\times K^{\prime}$. Furthermore, we define the mean and scale of the meta-modes respectively as $\overline{\boldsymbol{\mu}}_{\theta_{k^{\star}}}(X)=\frac{1}{K^{\prime}}\sum_{k^{\prime}}\pi_{\theta_{k^{\star},k^{\prime}}}(X)\boldsymbol{\mu}_{\theta_{k^{\star},k^{\prime}}}(X)$ and $\overline{\mathbf{b}}_{\theta_{k}^{\star}}^{2}(X)=\frac{1}{2K^{\prime}}\sum_{k^{\prime}}\left(2\mathbf{b}_{\theta_{k^{\star},k^{\prime}}}^{2}(X)+\right.$ $\left.\boldsymbol{\mu}_{\theta_{k^{\star},k^{\prime}}}^{2}(X)\right)-\overline{\boldsymbol{\mu}}_{\theta_{k^{\star}}}^{2}(X)$. Introducing these meta-modes leads to more stable predictions, as shown later in Tab. 3. We analyze further the loss robustness to the variation of the mode number in Appendix F. This hierarchical structure enables more spatial information as sub-modes can be expected to be close when sharing the same meta-mode. The following section describes the loss we developed to train an MDN with the concept of meta-modes.